The sum of 4 times Joan's age and 3 times Jim's age is 47. Jim is 1 year less than twice as old as Joan. Find their ages.
Answer by ptaylor(2198)

Let x=Joan's age; 4 times Joan's age=4x
Then 2x-1=Jim's age; 3 times Jim's age=3(2x-1)=6x-3
Now we are told that:
4x+6x-3=47 add 3 to each side
4x+6x-3+3=47+3 collect like terms
10x=50 divide each side by 10
x=5--------------------------------Joan's age
2x-1=2*5-1=9-----------------------Jim's age
CK
4*5 +3*9=47
20+27=47
47=47
Hope this helps---ptaylor